OpenStudy (anonymous):

PLEASE HELP!!! Will give a medal!!! What is the energy of a single photon of visible light with a wavelength of 500.0 nm? 1 nm = 10^ -9m

OpenStudy (anonymous):

I can partially answer your question. First we must find find the frequency of this wave. There is a relationship between wavelength and frequency, which is: \[\lambda v = c\] Where lampda is wavelength and 'c' is the speed of light. \[v = \frac{ c }{ \lambda }\] The energy is defined by this: \[E = hv = h \times \frac{ c }{ \lambda }\] Where 'h' is planck constant. I don't know the exact value of speed of light and the plack constant. That's all I remember. I hope it will help.

OpenStudy (anonymous):

Thanks, That does help!